Suppose cos(theta) = -5/13 and tan<0 find sin(2theta)
1
Expert's answer
2020-11-11T18:59:03-0500

"Cos\theta=\frac{-5}{13}"

"Cos^2\theta+sin^2\theta=1"

"(\frac{-5}{13})^2+sin^2\theta=1"

"Sin^2\theta=1-\frac{25}{169}"

"Sin^2\theta=\frac{144}{169}"

"Sin\theta=\frac{12}{13}"


"Sin2\theta=2sin\theta cos\theta"

"=2\times \frac{12}{13} \times \frac{-5}{13}"

"Sin2\theta=\frac{-120}{169}"